1. Let f(x) = √ x + 1 − 1 4−x2  . Compute the following: (10 points)

i. f(0)

ii. f(-3)

iii. f(2)

iv. f(-1)

v. f(3)

vi. Domain(f(x))


2. (a) Let f( x x−2 ) = 3x + 4 find f(x).

(b) Compute difference quotient of the function g(x) = √ x2 − 9, and simplify your answer.


3. Let f(x) = 1 − x, g(x) = x2 + bx + c. find b and c such that fog(x) = −x2  + 5x + 4. + bx + c. find b and c such that fog(x) = −x2+ 5x + 4.


1
Expert's answer
2021-01-29T12:45:51-0500

1. "f(x) = \sqrt{x+1} -14-x^2
\\ \\"

i) -13

ii) not exist

iii) "\sqrt{3}-18"

iv) -15

v) -21

vi) "[-1;+\infty)"

Explanation :

"\text { i) } f(0)=\sqrt{0+1}-14-0^{2}=1-14=-13 \\
\text { ii) } f(-3)=\sqrt{-3-1}-14-(-3)^{2}=\sqrt{-4}-14-9= \\
\text { - not defined) because } \sqrt{-4} \text { is not defined} \\
\text { iii) } f(2)=\sqrt{2+1}-14-(2)^{2}=\sqrt{3}-14-4=\sqrt{3}-18 \\
 \text { iv) } f(-1)=\sqrt{-1+1}-14-(-1)^{2}=0-14-1=-15 \\
\text { v) } f(3)=\sqrt{3+1}-19-3^{2}=2-14-9=-21 \\ 
\text { vi) } \sqrt{x+1} \text { needs } x+1 \geqslant 0 \text { , so domain is } x \in[-1 ;+\infty)"


2.

(a) "f(x^2-2) = 3x+4 \\"

Transform "x^2-2 \ to \ x"

"Let \ t=x^2-2, \quad so \ x = \pm\sqrt{t+2} \\
f(t) = 3x+4 \rightarrow f(x) = 3(\pm\sqrt{x+2}) + 4 = \pm3\sqrt{x+2} +4"

(b)

"If \ \ g(x) = \sqrt{x^2-9}
\\ Difference \ quotient
\\ \frac{g(x+h)-g(x)}{h}=\frac{\sqrt{(x+h)^{2}-9}-\sqrt{x^{2}-9}}{h} \\
If \ \ g(x) = \sqrt{x^2}-9 = |x| - 9 \\
\frac{g(x+h)-g(x)}{h}=\frac{|x+h|-9-|x|+9}{h}=\frac{|x+h|-|x|}{h}"

3.

"f(x)=1-x, g(x)=x^{2}+b x+c \\ 
(f \circ g)(x)=-x^{2}+5 x+4 \\"

"\begin{array}{l}
(f \circ g)(x)=f(g(x))=f\left(x^{2}+b x+c\right)=1-\left(x^{2}+b x+c\right)= \\
=1-x^{2}-b x-c=-x^{2}-bx+(1-c)=-x^{2}+5 x+4 \\
\Rightarrow-b=5,1-c=4 \Rightarrow b=-5, c=-3
\end{array}"